Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2000-051_2000-057

Mission Operations Days: 2000/051 to 2000/057
February 20, 2000 00:00:00 GMT to February 26, 2000 00:00:00 GMT


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
MODIS is in Science Mode - A Side
MODIS has no known Anomalies
ComponentState(s)Comment
BlackbodyA On (315 K); B OffNominal
Calibration ElectronicsA On; B OffNominal
Control ProcessorA On; B OffNominal
Nadir DoorUnlatched, OpenNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA On; B OffNominal
FDDI PortA On; B OffNominal
FIFO Memory#1 & #2 On; #3 & #4 OffNominal
Format ProcessorA On; B OffNominal
Power Supplies#1 On, #2 OffNominal
PV VIS FPAA On; B OffNominal
PV NIR FPAA On; B OffNominal
PV SMIR FPAA On; B OffNominal
PV LWIR FPAA On; B OffNominal
PC LWIR FPAA On; B OffNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOn (83 K)Nominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA On; B OffNominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA On; B OffNominal
Flight SoftwareRev BDNominal
Inhibit Ids SetNoneNominal
TMONs Enabled66, 67Nominal

This Week’s Completed MODIS Activities:

Saturday, February 19, 2000

2000/050 20:45:43 Real-time - - Set BB duty cycle to 100%
2000/050 21:12:22 ATC - - OA16: SD/SDSM Screened

Sunday, February 20, 2000

2000/051 03:46:55 ATC -- OA16: SD/SDSM Screened
2000/051 08:42:51 ATC -- OA16: SD/SDSM Screened
2000/051 14:00:00 ATC -- Sector shift to –3072, BB at 315K (EDS ~ 5 minutes)
2000/051 14:35:24 Real-time -- Set BB duty cycle to 33%
2000/051 14:36:18 Real-time – Set BB temp to 290K
2000/051 15:17:25 ATC -- OA15: SD/SDSM Open
2000/051 20:13:27 ATC -- OA16: SD/SDSM Screened (SSR buffer disables, no data)
2000/051 21:19 Real-time -- SMIR-A Vdet Sweeps, Itwk = 100, Vdet 207 – 231, by 1
2000/051 21:36 Real-time -- SMIR-A Vdet Sweeps, Itwk = 110, Vdet 207 – 231, by 1
2000/051 21:51 Real-time -- Reset SMIR-A default values, (Itwk = 79, VDET = 190)
2000/051 22:07 Real-time – Set Formatter to Test packets for 5 minutes

Monday, February 21, 2000

2000/052 02:48:27 ATC -- OA16: SD/SDSM Screened
2000/052 09:23:27 ATC -- OA15: SD/SDSM Open
2000/052 10:00 Bit Flip Investigation begins, MODIS OBC Activities suspended.
2000/052 10:24:00 Real-time -- Enable Formatter Test Data Packets (Bit-Flip Investigation)
2000/052 12:12 ATC -- Formatter Night -> Day -> Night -> Day -> Night (6 minutes) (Bit-Flip Investigation)
2000/052 16:35:32 Real-time – Return Formatter to Normal packets

Tuesday, February 22, 2000

2000/053 01:05 Bit Flip Investigation ends, MODIS OBC Activities resumed.
2000/053 03:29:42 ATC – OA16: SD/SDSM Screened
2000/053 07:30:38 ATC – Moved Formatter to Night up by 16:37 to support sector rotation
2000/053 08:20:00 ATC – Moved Formatter to Day up by 16:37 to support sector rotation
2000/053 08:25:47 – 08:37:35 ATC – Expeditied Data Service
2000/053 08:25:51 ATC – PV DCR Off
2000/053 08:25:53 ATC – PC DCR Off
2000/053 08:25:55 ATC – Sector Rotation to -3400
2000/053 08:25:57 ATC – OA16: SD/SDSM Screened
2000/053 08:37:27 ATC – Sector Rotation to 0
2000/053 08:37:29 ATC – PC DCR On
2000/053 08:37:31 ATC – PV DCR On
2000/053 09:09:23 ATC – Moved Formatter to Night up by 16:37 to support sector rotation
2000/053 13:22:12 ATC – OA15: SD/SDSM Open
2000/053 14:42:55 – 14:54:30 Real-time - - Ecal Sweeps (Part 1: NSTEP 20/20/20/8)
2000/053 14:54:48 – 15:45:19 Real-time - - Ecal Sweeps (Part 2: NSTEP 18/18/18/6)
2000/053 18:18:27 ATC - - OA16: SD/SDSM Screened
2000/053 19:29:50 – 19:41:37 Real-time - - Ecal Sweeps (Part 3: NSTEP 22/22/22/10)
2000/053 20:56 Real-time - - Vdet Sweeps (LWIR FPA Off)
2000/053 23:14:49 ATC - - OA16: SD/SDSM Screened

Wednesday, February 23, 2000

2000/054 04:11:23 ATC - - OA15: SD/SDSM Open
2000/054 09:07:57 ATC - - OA16: SD/SDSM Screened
2000/054 13:36:07 – 13:43:21 -- Real-time VDET sweeps(LWIR, Itwk=79)
2000/054 13:45:35 – 13:52:30 -- Real-time VDET sweeps(LWIR, Itwk=85)
2000/054 14:04:31 ATC - - OA16: SD/SDSM Screened
2000/054 14:26:50 – 14:33:45 -- Real-time VDET sweeps(LWIR, Itwk=90)
2000/054 14:36:06 – 14:41:16 -- Real-time VDET sweeps(LWIR, Itwk=95)
2000/054 14:41:46 – 14:42:36 -- Real-time VDET back to nominal (Itwk=79)
2000/054 15:20:28 – 15:21:22 -- Real-time SMIR-B turned on
2000/054 15:21:49 - 15:28:55 -- Real-time VDET sweeps(SMIR-B, Itwk=79)
2000/054 15:30:52 – 15:37:47 -- Real-time VDET sweeps(SMIR-B, Itwk=90)
2000/054 15:55:13 – 16:02:07 -- Real-time VDET sweeps(SMIR-B, Itwk=100)
2000/054 16:04:43 – 16:11:28 -- Real-time VDET sweeps(SMIR-B, Itwk=110)
2000/054 16:11:44 – 16:12:24 -- Real-time VDET back to nominal (Itwk=79)
2000/054 16:12:38 – 16:13:32 -- Real-time SMIR-A turned on
2000/054 19:01:05 ATC - - OA15: SD/SDSM Open

Thursday, February 24, 2000

2000/055 01:33:20 ATC – OA16: SD/SDSM Screened Double
2000/055 08:08:46 ATC – OA16: SD/SDSM Screened Double
2000/055 14:44:11 ATC – OA15: SD/SDSM Open Double
2000/055 15:27:00 Real-time - - OPEN NADIR APERTURE DOOR (Earth View Data!)
2000/055 15:30:00 Real-time - - Dump door table
2000/055 20:04:45 – 20:13:14 Real-time - - SMIR VDET sweeps (Itwk=125)
2000/055 20:15:17 – 20:23:18 Real-time - - SMIR VDET sweeps (Itwk=155)
2000/055 20:23:41 – 20:31:37 Real-time - - SMIR VDET sweeps (Itwk=140)
2000/055 20:32:04 – 20:40:03 Real-time - - SMIR VDET sweeps (Itwk=170)
2000/055 20:40:19 – 20:40:59 Real-time - - Set SMIR VDET back to nominal (Itwk=79)
2000/055 21:19:38 ATC – OA16: SD/SDSM Screened Double

Friday, February 25, 2000

2000/056 03:55:32 ATC – OA16: SD/SDSM Screened Double
2000/056 10:31:03 ATC – OA16: SD/SDSM Open Double
2000/056 17:06:34 ATC – OA16: SD/SDSM Screened Double
2000/056 18:39:33 Real-time - - Set BB Temp to 315K
2000/056 18:40:29 Real-time - - Set BB duty cycle to 100%
2000/056 22:00 Real-time - - Sector Rotation to –3072 for 5 minutes for viewing BB.
2000/056 22:05 Real-time - - Sector Rotation to 0 home position.
2000/056 22:06 Real-time - - Set BB Temp to 270K
2000/056 22:42 Real-time - - Turn off BB
2000/056 23:42:05 ATC – OA16: SD/SDSM Screened Double

Saturday, February 26, 2000

2000/057 06:17:36 ATC – OA16: SD/SDSM Open Double
2000/057 12:53:07 ATC – OA16: SD/SDSM Screened Double
2000/057 17:05 Real-time - - Set Vdet/Itwk to new nominal settings (TBD)
2000/057 17:06 Real-time - - Set BB temp to 315K
2000/057 18:20:29 – 18:28:31 Real-time - - Repeat Vdet Sweeps (Itwk 125, BB=298K)
2000/057 18:30:41 – 18:38:33 Real-time - - Repeat Vdet Sweeps (Itwk 155, BB=298K)
2000/057 18:38:54 – 18:39:37 Real-time - - Reset Itwk/Vdet to nominal (110/226)
2000/057 19:28:39 ATC – OA16: SD/SDSM Screened Double

This Week's Scheduled MODIS Activities Not Completed:

None.

Upcoming MODIS Events:

Saturday, February 26th, 2000

2000/057 20:30 Real-time - - Turn off BB

Sunday, February 27th, 2000

2000/058 02:04:11 ATC – OA15: SD/SDSM Open Double
2000/058 08:39:42 ATC – OA16: SD/SDSM Screened Double
2000/058 14:30(?) Real-time - - Set Itwk/Vdet to 100/220
2000/058 14:30(?) Real-time - - Tun on BB and set to 315K
2000/058 15:15:13 ATC – OA16: SD/SDSM Screened Double
2000/058 18:00(?) Real-time - - Turn off BB
2000/058 21:50:44 ATC – OA15: SD/SDSM Open Double

Monday, February 28th, 2000

2000/059 04:26:15 ATC – OA16: SD/SDSM Screened Double
2000/059 11:01:46 ATC – OA15: SD/SDSM Open Double
2000/059 12:00:00 ATC – SRCA Full Radiometric
2000/059 12:40:39 ATC – OA16: SD/SDSM Screened Double
2000/059 14:00(?) Real-time - - Reset Itwk/Vdet to nominal values (110/226)
2000/059 14:00(?) Real-time - - Turn on BB and set to 290K
2000/059 15:00(?) Real-time - - Set BB duty cycle to 33%
2000/059 22:33:56 ATC – OA16: SD/SDSM Screened Double

Tuesday, February 29th, 2000

2000/060 03:30:34 ATC – OA16: SD/SDSM Screened Double
2000/060 08:27:12 ATC – OA15: SD/SDSM Open Double
2000/060 12:00:00 ATC – SRCA Full Spatial
2000/060 15:02:43 ATC – OA16: SD/SDSM Screened Double
2000/060 19:59:22 ATC – OA16: SD/SDSM Screened Double

Wednesday, March 1, 2000

2000/061 06:44:57 ATC – SRCA Full Spectral, 30W part I
2000/061 08:39:41 ATC – SRCA Full Spectral, 30W part II
2000/061 10:02:06 ATC – SRCA Full Spectral, 10W part I
2000/061 11:57:26 ATC – SRCA Full Spectral, 10W part II

MODIS Anomalies:

At 2000/051 08:21:25 MOD_VR_PVSM_P5VD6_12 exceeded its yellow low limit with a value of 4.61 for one 65 second cycle.

General Instrument Comments:

MODIS is in Science Mode on the A-side with the SVD and NAD open. Functional checkout of MODIS is complete and was successful. All mechanisms behaved nominally.

First light occurred at 2000/055 15:27:09

MODIS Telemetry Trends:

Telemetry is nominal.

Non-MODIS Significant Events:

MISR opened their covers on 2000/055 and all went nominally.
CERES successfully opened both of their covers on 2000/056.
MOPITT is schedule to open their covers on Monday, 2000/059.

The following is the as-run burn schedule which placed Terra into her final orbit (the times are in GMT). Our current orbit is 708km x 692km. FDS did such a great job that another orbit attainment burn will not be required for about 2 months (around the time of the inclination maneuver). Therefore, we’ll be skipping the March monthly orbit attainment burn. Kudos to FDS.

Burn | Start Time | End Time | Duration(sec)

1 | Feb 10 2000 21:15:43.030 | Feb 10 2000 21:16:43.030 | 60.000 (complete)
2 | Feb 12 2000 20:43:22.876 | Feb 12 2000 20:45:52.876 | 150.000 (complete)
3 | Feb 14 2000 20:12:16.920 | Feb 14 2000 20:17:16.920 | 300.000 (complete)
4 | Feb 16 2000 19:45:32.361 | Feb 16 2000 19:50:52.361 | 320.000 (complete)
5 | Feb 18 2000 21:52:11.699 | Feb 18 2000 21:57:31.699 | 320.000 (complete)
6 | Feb 20 2000 19:03:07.853 | Feb 20 2000 19:08:27.853 | 320.000 (complete)
7 | Feb 22 2000 18:00:01.007 | Feb 22 2000 18:04:41.107 | 280.000 (complete)
8 | Feb 23 2000 21:10:21.094 | Feb 24 2000 21:12:11.094 | 110.000 (complete)

Getting data through the system and products delivered to MCST continues to be a major problem and is the critical path for characterization of MODIS.

EDOS has installed their bit flip patch Friday evening (2000/057) and initial indications are that this has greatly increased EDOS ability to process MODIS data. Thanks EDOS!

Limited Life Item Status:

All limited life items are well within lifetime ranges. The precise statistics for each item have been received from LMMS/Valley Forge and will be incorporated next week.